OpenStudy (anonymous):

Based on the data set shown, which of the following is a true statement? -1, -1, 0, 1, 1, 1, 1, 2, 2, 2,3 mean = mode mean < median mode < median

OpenStudy (anonymous):

What about the 3?

OpenStudy (ajprincess):

As u see mean is equal to median 3 cannt b the choice.

OpenStudy (anonymous):

No I was saying that you left out the 3

OpenStudy (ajprincess):

oops sorry.

OpenStudy (anonymous):

Its finnee

OpenStudy (ajprincess):

Mean=\(\frac{(-1)+(-1)+0+1+1+1+1+2+2+2+3}{11}\) =\(\frac{11}{11}\) =1 Mode=1 Median=\(\frac{N+1}{2}\)th term =\(\frac{11+1}{2}\)th term =\(\frac{12}{2}\)th term =6th term =1 Nw can u do?

OpenStudy (ajprincess):

This time I am pretty sure I am right.

OpenStudy (anonymous):

So everything =1? And How come tthe mode is not 1,-1, and 2?

OpenStudy (ajprincess):

mode is the number that comes many times in a set of data. -1 comes two times. 0 comes one time 1 comes four times 2 comes three times 3 comes one time I thnk nw u will be able to decide which is mode.

OpenStudy (anonymous):

Yup thanks. So the Mean = the Mode right?

OpenStudy (ajprincess):

yes.

OpenStudy (anonymous):

Thanks again

OpenStudy (ajprincess):

yw.:)

OpenStudy (anonymous):

Lol cann i have a medal?

OpenStudy (ajprincess):

oh sure. sorry for nt givng at first.

OpenStudy (anonymous):

Oh its fine lol. I just like medals